OpenStudy (anonymous):

What is the equation of a vertical line passing through the point (–5, –1)? y = –5 y = –1 x = –5 x = –1

OpenStudy (anonymous):

x = -5

OpenStudy (bahrom7893):

any vertical line means x = __

OpenStudy (anonymous):

y varies directly as the square of x. When x = 3, y = 54. Find y when x = 8.

OpenStudy (bahrom7893):

y = 6x^2

OpenStudy (anonymous):

y = 96 y = 144 y = 256 y = 384

OpenStudy (bahrom7893):

y = 6*64 = 384

OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8)?

OpenStudy (anonymous):

y = 2(x + 2) y = 2(x – 2) y = –2(x – 2) y = –2(x + 2)

OpenStudy (bahrom7893):

Slope: (8-0)/(2+2) = 2

OpenStudy (bahrom7893):

y - 0 = 2(x-(-2)) y = 2(x+2)

OpenStudy (anonymous):

What is the equation of the line passing through (–3, 6) and parallel to the line y = 4x – 2 in standard form?

OpenStudy (anonymous):

x – 4y = –18 x + 4y = 18 4x – y = –18 4x + y = 18